sam is 34yrs older than ben 7yrs from now sam's age will be 3 times as old as ben how old are they now?

Answer by ankor@dixie-net.com(22740)

sam is 34yrs older than ben
s = b + 34
7yrs from now sam's age will be 3 times as old as ben
s + 7 = 3(b+7)
s + 7 = 3b + 21
s = 3b + 21 - 7
s = 3b + 14
Replace s with (b+34) and solve
how old are they now?